3 times a number, increased by 7, is 19. Find the number.
Answer by doukungfoo(195)

Transforming an english sentence into an algebra equation is just a matter of breaking the sentence down and identifying how each part relates to an algebra equation.
Lets call the number x
The first part of the sentence reads "3 times a number"
times means multiply so the first part of our equation is
3x
The next part tells us "increased by 7"
if you increase something you are adding to it
so now we have
3x + 7
The last part of the sentence says "is 19"
The word "is" tells us we need an equal sign
so now we have the complete equation
3x + 7 = 19
Lets solve for x
first subtract 7 from both sides
3x + 7 - 7 = 19 - 7
3x = 12
next divide both sides by 3
3x/3 = 12/3
x = 4
Answer: The number is 4
check the answer by replacing x with 4
3x + 7 = 19
3(4) + 7 = 19
12 + 7 = 19
19 = 19